# Streamlined Federated Unlearning: Unite as One to Be Highly Efficient

### v1

@@ V-A2 Training Settings
V-A3 Baseline Methods
In addition to comparing the most basic method of retraining, we also compared it with the SOTA (state-of-the-art) method FUCDP for unlearning class data in FU.
V-B Results and Analysis
The evaluation of unlearning methods typically considers three key dimensions: the effectiveness of unlearning the target data (unlearning efficacy), the preservation of model performance on the retained data after unlearning (unlearning fidelity), and the efficiency in terms of time and communication (unlearning efficiency). We conducted extensive experiments, with selected results presented. It is important to note that FUCDP is only applicable to CNN models, while we employed an LSTM model for training on the DBpedia dataset, demonstrating the broader applicability of SFU.
@@ V-B Results and Analysis
Figure 3: Test accuracy of models using different unlearning methods on forgotten and retained data. We unlearned three different classes of data on each of the three different datasets using different methods. These category data are randomly distributed on different clients according to non-iid. The lines of different colors represent the unlearning effects of different methods. The numbers of different colors on the horizontal axis represent the number of communication rounds required by each method to complete unlearning.
@@
Figure 4: Comparison of the unlearning performance of SFU with different hyperparameters $\alpha$. We set three different orders of magnitude of $\alpha$ to compare its impact on unlearning: $\alpha$ calculated according to the proposed method and $\alpha$ one order of magnitude above and below it. We select one class of data from each of the three datasets for unlearning.
V-B1 Unlearning Efficacy
We first compared the prediction accuracy of the models on the forgotten data before and after unlearning. Theoretically, the more thoroughly the data is unlearned, the closer this value should be to that of retraining, ideally approaching zero. As shown in Table I, we randomly perform class data unlearning for each dataset and present the unlearning results for three selected classes. The results indicate that both SFU and FUCDP demonstrate results on the image datasets CIFAR-10 and CIFAR-100 that are very close to retraining. Similarly, on the text dataset Dbpedia, SFU achieves results comparable to retraining, with accuracy nearing zero.
[…]
V-B2 Unlearning Fidelity
We assess unlearning fidelity by comparing the prediction accuracies of the models on the retained data before and after unlearning. As shown in Table I, both SFU and FUCDP achieve accuracies comparable to those before unlearning, indicating minimal or no degradation in model performance. However, they achieve this in different ways. As illustrated in Figure 3, FUCDP undergoes a performance recovery step, utilizing a fine-tuning process akin to retraining and requiring a certain number of rounds to restore performance to its previous state before influence removal. In contrast, SFU maintains model performance consistently around the pre-unlearning state, thereby eliminating the need for a performance recovery step.
V-B3 Unlearnng Efficiency
With unlearning efficacy and fidelity ensured, the most crucial metric for evaluating an FU method is unlearning efficiency, which is the primary motivation behind the development of unlearning techniques. We use two parameters for comparison. Considering the communication cost in the FL environment, the first parameter is the comparison of communication rounds across different approaches. The second parameter is the speedup ratio in terms of time compared to the baseline approach (retraining).
As shown in Table II, while FUCDP demonstrates a significant improvement over retraining, SFU surpasses FUCDP in both time efficiency and communication rounds. Particularly, SFU completes unlearning on all three datasets in just 1-2 communication rounds on average, significantly reducing communication costs. Regarding the time acceleration ratio, SFU shows a more pronounced speedup on CIFAR-10 compared to CIFAR-100 and Dbpedia, which can be attributed to the higher complexity of CIFAR-100 and Dbpedia. Overall, SFU significantly enhances unlearning efficiency compared to existing methods.
@@ V-B3 Unlearnng Efficiency
TABLE III: Performance of SFU when unlearning multiple classes of data. We show the performance of SFU for the first class data and the last class data of unlearning, and also the overall time speedup ratio and the number of communication rounds required.
@@
Figure 5: The unlearning performance of SFU under different combinations of teacher models.
V-B4 Hyperparameters
Given the varying amounts of target data and residual data across different datasets and unlearning requests, we introduce a hyperparameter in the loss function to balance these two aspects. From the perspective of the teacher models, we aim for a balance between ”knowledge of forgetting” and ”knowledge of performance” when guiding the student model. This balance is crucial to effectively removing the influence of the target data without compromising overall model performance.
To accommodate different datasets, we define the parameter $\alpha$ as $\alpha={\text{len}(x^{r})}/{\text{len}(x^{f})}$, where $x^{r}$ represents the retained data and $x^{f}$ the target data to be forgotten. We conducted extensive experiments to validate the effectiveness of this parameter across various values of $\alpha$. The results, presented in Figure 4, indicate that the choice of $\alpha$ significantly influences the balance between performance retention and influence removal. For instance, in the CIFAR-10 dataset when unlearning class 5, we compared the default $\alpha=9$ (as defined) with $\alpha=1$ and $\alpha=50$. The results show that $\alpha=1$ preserves model performance but fails to effectively remove the target data’s influence, while $\alpha=50$ successfully removes the influence but at the cost of decreased performance. The default $\alpha$ achieves a balance between these extremes, maintaining performance while effectively removing the influence of forgotten data. Similar observations were made for CIFAR-100 and the text dataset Dbpedia. In conclusion, when $\alpha$ is too small, the model retains performance but fails to remove the target data’s influence effectively. Conversely, a large $\alpha$ results in effective influence removal but degrades performance. Our definition of $\alpha$ thus strikes an optimal balance, ensuring that the model removes the influence of forgotten data without compromising its performance and is adaptable to various datasets and unlearning requests.
V-B5 Multiple classes unlearning
For multiple classes unlearning, SFU only needs to loop through the execution after completing a single class. The results are shown in Table III, where we run multiple classes unlearning experiments on image and text datasets respectively. It is obvious that the efficiency of SFU unlearning is about the same as that of single-class unlearning. At the completion of the last class unlearning, the accuracy of the model on the retained data is still very close to the state before unlearning, which also shows that the model performance is well preserved. Moreover, at the completion of unlearning for each class, the model’s accuracy on the forgotten data was close to 0, similar to that of retraining, which also indicates that the model did a good job of removing the influence of the target data. Most importantly, for each class that needs to be forgotten, SFU only needs 1 or 2 rounds of communication to complete the unlearning, which greatly improves communication efficiency and time efficiency.
V-C Ablation study
@@
We developed a multi-teacher system comprising three distinct teacher models: the performance preservation teacher model ($M^{P}$), the forgetting teacher model ($M^{F}$), and the label-based preservation teacher model ($M_{i}^{L}$). Based on their roles, $M^{F}$ is responsible for forgetting, while $M^{P}$ and $M^{L}$ focus on performance preservation. To validate the rationale behind this design and assess the necessity of each teacher model, we conducted extensive ablation experiments. Specifically, we performed three sets of comparative experiments: using only the $M^{P}$ and $M^{F}$ combination, only the $M^{L}$ and $M^{F}$ combination, and all three models ($M^{P}$, $M^{L}$, and $M^{F}$) together. As shown in Figure 5, when unlearning class 5 in the CIFAR-10 dataset, all three combinations effectively removed the influence of the forgotten data. However, in terms of performance preservation, the combinations of $M^{P}$ with $M^{F}$ or $M^{L}$ with $M^{F}$ led to some performance degradation. Only the full combination of $M^{P}$, $M^{L}$, and $M^{F}$ successfully removed the influence of the forgotten data while maintaining model performance, thereby significantly enhancing unlearning efficiency. In summary, each teacher model in our multi-teacher system serves a unique and essential role, making them all indispensable to the system’s overall effectiveness.
VI Conclusion
@@
In this study, we propose a streamlined federated unlearning (SFU) method that significantly enhances unlearning efficiency by integrating the steps of influence removal and performance recovery. We designed a multi-teacher system where three distinct teacher models guide the unlearned model to remove the influence of target data while maintaining model performance. SFU does not require additional computational or storage resources. It adheres to the security requirements of federated learning (FL) without necessitating global data access and is not constrained by data distribution. Furthermore, SFU exhibits great flexibility, seamlessly integrating into any training stage of FL. Experimental results demonstrate that SFU is highly generalizable, applicable to various models, and effective on both text and image datasets. Compared to baseline retraining methods, SFU significantly improves time and communication efficiency and substantially outperforms existing SOTA methods. Additionally, the effectiveness of SFU is validated through backdoor attacks. We believe this work greatly enhances the privacy of federated learning and improves the efficiency of federated unlearning. In future work, we will continue to explore more efficient and general federated unlearning methods.

### v2

@@ VI-A2 Training Settings
VI-A3 Baseline Methods
In addition to comparing the basic method of retraining, we also compare it with the SOTA methods FUCDP [36] and FUR [16] for unlearning class data in FU. FUCDP achieves unlearning by computing the TF-IDF scores of different category channels and pruning the channels with high scores for the target forgotten category. While pruning enables fast unlearning, it also significantly degrades model performance, necessitating fine-tuning for recovery. Similarly, FUR removes specific gradient residues and adds Gaussian noise to achieve the unlearning of specific category data. However, this method also leads to a decline in model performance. Although researchers claim that this degradation is negligible, in reality, the impact remains substantial. Therefore, we introduce an additional fine-tuning step in the unlearning process of FUR to restore model performance.
VI-B Results and Analysis
The evaluation of unlearning methods is typically conducted in three dimensions: unlearning efficacy, unlearning fidelity, and unlearning efficiency. Unlearning efficacy assesses the extent to which the forgotten data has been successfully unlearned, indicating whether the unlearning process has been fully completed. Unlearning fidelity is used to describe the performance of the model on the retained data after the unlearning process. A good FU method must ensure the continued availability of the model even after unlearning has been completed. Unlearning efficiency refers to the time and communication efficiency of the FU method in completing unlearning. Once the first two evaluation perspectives have been completed, this dimension provides the most direct means of measuring an FU method, as it reflects the fundamental objective of FU. We conduct extensive experiments, with selected results presented. It is important to note that FUCDP is only applicable to CNN models, and FUR is difficult to extend to text datasets. Therefore, we employ an LSTM model for training on the DBpedia dataset, demonstrating the broader applicability of SFU.
@@ VI-B Results and Analysis
Figure 3: Test accuracy of models using different unlearning methods on forgotten and retained data. Three different classes of data on each of the three different datasets are randomly selected to be unlearned using different methods. These data are randomly distributed on different clients according to non-iid. The lines of different colors represent the unlearning effects of different methods. The numbers of different colors on the horizontal axis represent the number of communication rounds required by each method to complete unlearning.
VI-B1 Unlearning Efficacy
We employ two metrics to evaluate unlearning efficacy: the accuracy of the model on forgotten data before and after unlearning, and the success rate of backdoor attacks. Firstly, we compare the prediction accuracy on forgotten data before and after unlearning across different FU methods. Theoretically, the more thoroughly the model unlearns, the closer this value approaches the prediction accuracy of retraining, which is nearly 0. As shown in Table III, we randomly select class data from each dataset for unlearning and report results for three specific classes. For the image datasets CIFAR-10 and CIFAR-100, SFU, FUCDP, and FUR achieve post-unlearning prediction accuracies that are nearly identical to the retraining baseline, approaching 0. On the additional text dataset Dbpedia, SFU similarly attains a prediction accuracy close to that of retraining, also approaching 0. From the perspective of prediction accuracy, these methods demonstrate unlearning efficacy comparable to retraining, effectively achieving comprehensive removal of forgotten data.
@@ VI-B1 Unlearning Efficacy
VI-B2 Unlearning Fidelity
We evaluate unlearning fidelity by examining the prediction accuracy of the model on retained data before and after unlearning. Ideally, an effective unlearning process should only affect forgotten data while preserving the model performance on retained data. As shown in Table III, when unlearning is performed using retraining, the prediction accuracy on retained data remains almost identical to its pre-unlearning accuracy. Similarly, after unlearning with SFU and FUCDP, the prediction accuracy on retained data remains close to its original level. However, for FUR, the results presented in the table show a noticeable decline in model performance after unlearning, which is non-negligible. This indicates that FUR negatively impacts retained data during the unlearning process. To address this issue, we apply additional fine-tuning to restore model performance. In contrast, FUCDP inherently includes a fine-tuning step, which allows it to recover performance effectively, as observed in the table.
The detailed process is illustrated in Figure 3, where the prediction accuracy on retained data changes over different training iterations. When unlearning is performed using retraining (green line), the accuracy gradually increases. FUCDP (blue line) starts with pruning and then uses fine-tuning to restore model performance. The figure shows the process of FUCDP fine-tuning, where the prediction accuracy on the retained data slowly increases. Similarly, FUR (brown line) exhibits a recovery trend after fine-tuning, eventually restoring prediction accuracy to pre-unlearning levels. For better visualization, the unlearning epochs for SFU are extended. The results clearly show that SFU (red line) maintains a stable prediction accuracy on retained data throughout the unlearning process, without any noticeable degradation. This suggests that SFU effectively preserves model performance during unlearning, eliminating the need for additional fine-tuning.
On the other hand, the prediction accuracy on the forgotten data rapidly decreases for all methods. The retraining approach (green dashed line) maintains an accuracy of 0, while FUCDP (blue dashed line) and FUR (yellow dashed line) stabilize near 0 after unlearning. SFU (red dashed line) also rapidly drops to 0 within very few iterations. This trend aligns with the conclusions from the previous section, confirming that all methods successfully remove information from forgotten data while exhibiting varying effects on retained data. Additionally, the numbers on the x-axis in Figure 3 represent the number of training epochs required for each FU method to complete unlearning. It is evident that the main factor influencing efficiency is the recovery of model performance. FUCDP and FUR require additional fine-tuning steps to restore accuracy, following a retraining-like process that demands more iterations. In contrast, SFU maintains stable model performance throughout unlearning, eliminating the need for fine-tuning and demonstrating superior efficiency and fidelity.
VI-B3 Unlearnng Efficiency
We use two metrics to measure the unlearning efficiency of these FU methods: number of communication rounds and speed-up ratio in time compared to retraining. As shown in Table IV, in terms of communication rounds, the results across all datasets indicate that SFU, FUR, and FUCDP require significantly fewer communication rounds to complete unlearning than retraining. Among them, SFU demonstrates the greatest efficiency, requiring only 1-2 rounds on average, substantially reducing communication overhead and outperforming both FUR and FUCDP.
Regarding the speedup ratio, while FUCDP and FUR already achieve significant improvements over retraining, SFU consistently exhibits a much higher speedup across all datasets, highlighting its superior computational efficiency. This suggests that SFU not only ensures effective unlearning but also significantly reduces computational costs, making it a more efficient approach. Additionally, we observe that the reduction in communication rounds and the improvement in speedup ratio are more pronounced on CIFAR-10 and DBpedia compared to CIFAR-100. This is likely because CIFAR-100 has a larger number of classes and higher task complexity, making unlearning more challenging. Similarly, while DBpedia is a text dataset, its higher information density increases processing difficulty compared to CIFAR-10, leading to variations in efficiency across datasets. In conclusion, SFU significantly enhances unlearning efficiency for class data, far surpassing the baseline retraining method and outperforming the current SOTA methods FUCDP and FUR.
@@ VI-B3 Unlearnng Efficiency
TABLE V: Performance of SFU when unlearning multiple classes of data. We show the performance of SFU for the first-class data and the last-class data of unlearning, as well as the overall time speedup ratio and the number of communication rounds required.
@@ VI-C Ablation Study
Figure 4: The unlearning performance of SFU under different combinations of teacher models. The lines of different colors represent the unlearning effects of different combinations of teacher models
[…]
Figure 5: Comparison of the unlearning performance of SFU with different values of hyperparameters $\alpha$. We set three different orders of magnitude of $\alpha$ to compare its impact on unlearning: $\alpha$ calculated according to the proposed method and $\alpha$ one order of magnitude above and below it. We select one class of data from each of the three datasets for unlearning.
VI-C1 Combinations of Teachers Models
@@
We develop a multi-teacher system comprising three distinct teacher models: the performance preservation teacher model ($M^{P}$), the forgetting teacher model ($M^{F}$), and the label-based preservation teacher model ($M_{i}^{L}$). Based on their roles, $M^{F}$ is responsible for forgetting, while $M^{P}$ and $M^{L}$ focus on performance preservation. To validate the rationale behind this design and assess the necessity of each teacher model, we conduct extensive ablation experiments. Specifically, we perform three sets of comparative experiments: using only the $M^{P}$ and $M^{F}$ combination, only the $M^{L}$ and $M^{F}$ combination, and all three models ($M^{P}$, $M^{L}$, and $M^{F}$) together. As shown in Figure 4, all three combinations effectively remove the influence of forgotten data on three datasets. However, in terms of performance preservation, the combinations of $M^{P}$ with $M^{F}$ or $M^{L}$ with $M^{F}$ led to some performance degradation. Only the complete combination of $M^{P}$, $M^{L}$, and $M^{F}$ successfully remove the influence of forgotten data while maintaining model performance, thus significantly enhancing the efficiency of unlearning. In summary, each teacher model in our multi-teacher system serves a unique and essential role, making them all indispensable to the overall effectiveness of the system.
VI-C2 Hyperparameters
Given the varying amounts of target data and residual data across different datasets and unlearning requests, we introduce a hyperparameter in the loss function to balance these two aspects. From the perspective of the teacher models, we aim for a balance between “knowledge of forgetting” and “knowledge of performance” when guiding the student model. This balance is crucial to effectively removing the influence of the target data without compromising overall model performance.
To accommodate different datasets, we define $\alpha={num(x^{r})}/{num(x^{f})}$, where $x^{r}$ represents the retained data and $x^{f}$ the target data to be forgotten. We conduct extensive experiments to validate the effectiveness of this parameter across various values of $\alpha$. The results, presented in Figure 5, indicate that the choice of $\alpha$ significantly influences the balance between performance retention and influence removal. For instance, in the CIFAR-10 dataset when unlearning class 5, we compare the default $\alpha=9$ (as defined) with $\alpha=1$ and $\alpha=50$. The results show that $\alpha=1$ preserves model performance but fails to effectively remove the influence of the target data, while $\alpha=50$ successfully removes the influence but at the cost of decreased performance. The default $\alpha$ achieves a balance between these extremes, maintaining performance while effectively removing the influence of forgotten data. Similar observations were made for CIFAR-100 and the text dataset Dbpedia. In conclusion, when $\alpha$ is too small, the model retains performance but fails to remove the influence of the forgotten data effectively. Conversely, a large $\alpha$ results in effective influence removal but degrades performance. Our definition of $\alpha$ thus strikes an optimal balance, ensuring that the model removes the influence of forgotten data without compromising its performance and is adaptable to various datasets and unlearning requests.
VI-C3 Multiple classes unlearning
For multiple classes unlearning, SFU only needs to loop through the execution after completing a single class. The results are shown in Table V, where we run multiple classes unlearning experiments on image and text datasets, respectively. It is obvious that the efficiency of SFU unlearning is about the same as that of single-class unlearning. At the completion of the last class unlearning, the accuracy of the model on the retained data is still very close to the state before unlearning, which also shows that the model performance is well preserved. Moreover, at the completion of unlearning for each class, the accuracy on the forgotten data was close to 0, similar to that of retraining, which also indicates that the model did a good job of removing the influence of the target data. Most importantly, for each class that needs to be forgotten, SFU only needs 1 or 2 rounds of communication to complete the unlearning, which greatly improves communication efficiency and time efficiency.
VII Conclusion
@@
In this study, we propose a streamlined federated unlearning method called SFU that significantly enhances unlearning efficiency by integrating the steps of influence removal and performance recovery. We design a multi-teacher system where three distinct teacher models guide the unlearned model to remove the influence of target data while maintaining model performance. SFU does not require additional computational or storage resources. It adheres to the security requirements of FL without necessitating global data access and is not constrained by data distribution. Furthermore, SFU exhibits great flexibility, seamlessly integrating into any training stage of FL. Experimental results demonstrate that SFU is highly generalizable, applicable to various models, and effective on both text and image datasets. Compared to retraining, SFU significantly improves time and communication efficiency and substantially outperforms existing SOTA methods. Additionally, the effectiveness of SFU is validated through backdoor attacks. In the future, we plan to explore more efficient and generalized FU methods for more unlearning scenarios and goals.